ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 08, 22, 54, 40 Choban, Mitrofan M., Tiraspol State University, Moldova Algebraical equivalences of topological spaces We present the general aspect of algebraical equivalences of topological spaces. The M-equivalence and A-equivalence of topological spaces are particular cases of this general conception. We prove that in larger classes of spaces the covering dimension, paracompactness, p-paracompactness, lindelöfness and metrizability are invariant properties of all algebraical equivalences. References: 1. A. V. Arhangel’skii, Classes of topological groups, Russian Math. Surveys 36:3 (1981), 151–174; 2. M. M. Choban, Some topics in topological algebra, Topol. Appl. 54 (1993), 183–202; 3. M. I. Graev, Free topological groups, Trans. AMS 8 (1962), 303–364.